Alexandra brought a hat that costs 14 dollars and a watch that costs 6 dollars. At the counter she received a discount. If she payed only 15 dollars, how much was the discount?
1
Expert's answer
2020-04-22T07:35:04-0400

cost of hat = $ 14

cost of watch = $ 6


Total cost = 14+6 = $20

She paid at counter = $ 15


Discount she received = Total cost - Paid at counter

= $ 20 - $ 15

= $ 5

So discount she received is $ 5

In percentage

discount percentage = (discount/total cost)*100

=(5/ 20)*100

=0.25*100

= 25 %

so she receives 25% as discount